# Geometric Rectification of Surface Activity Induced Flow in Confined Channels

### v1

@@ Abstract
Conventional pressure-driven flow obeys Poiseuille’s law, with the mean velocity scaling as $u\propto r^{2}$ under confinement. Here we identify a distinct transport mode driven by spatially structured surface activity (e.g., mass exchange or boundary slip gradients), which is rectified by geometric asymmetry into a net axial flux. Using a minimal exchange model, we show that this mechanism exhibits four defining signatures that are inconsistent with classical lubrication theory: (i) an inverted confinement scaling $u\propto r^{-1}$ (“narrower-is-faster”); (ii) leading-order viscosity independence; (iii) macroscopic length amplification ($Q\propto L$); and (iv) linear superposition with pressure-driven flows. These results establish confined channels as active geometric rectifiers and provide a unified framework for surface-induced transport from microfluidic to biological settings.
@@
Using the reciprocal theorem, the pressure-free throughput can be written as the projection
$Q=-\int_{0}^{L}s(x)\,\phi(x)\,\mathrm{d}x.$

### v2

@@ Abstract
Conventional pressure-driven flow obeys Poiseuille’s law, with the mean velocity scaling as $u\propto r^{2}$ under confinement. Here we identify a distinct transport mode driven by surface activity (e.g., mass exchange or boundary slip gradients), which is rectified by geometric asymmetry into a net axial flux. Using a minimal exchange model, we show that this mechanism exhibits four defining signatures that contrast sharply with pressure-driven Poiseuille flow: (i) an inverted confinement scaling $u\propto r^{-1}$; (ii) leading-order viscosity independence for prescribed $s$; (iii) macroscopic length amplification ($Q\propto L$); and (iv) linear superposition with pressure-driven flows. These results establish confined channels as active geometric rectifiers and provide a unified framework for surface-induced transport from microfluidic to biological settings.